OpenStudy (anonymous):

how do I solve this using perfect square 3x^2-6x=17 I need the answer with work

OpenStudy (anonymous):

are you familiar with completing the square?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

am I wrong or you'd need the quadratic formula for that?

OpenStudy (anonymous):

I need both

OpenStudy (anonymous):

ax^2 + bx + c = 0 turn that into: a(x+d)^2 + e = 0 where d = b/(2a) and e = c - (b^2)/(4a)

OpenStudy (jdoe0001):

\( 3x^2-6x-17=0 \), and plug the values in the quadratic formula :|

OpenStudy (anonymous):

that doesn't allow him to solve for perfect squares

OpenStudy (jdoe0001):

ok, I see the difference